# Entanglement classification and non-k-separability certification via Greenberger-Horne-Zeilinger-class fidelity

### v1

@@ I Introduction
One of the important consequences of the quantum description of physical systems is the notion of entanglement [horodecki2009entanglement] which is the necessary resource for the Einstein-Podolsky-Rosen steering [uola2020steering] and the Bell correlations [bell, brunner2014bell, bell_rmp]. None of these correlations has a classical counterpart. The highlighted role of entanglement requires a systematic characterisation and classification of quantum states in terms of some entanglement measure [PhysRevLett.78.2275, Horodecki2001, Plenio_2007]. Nevertheless, the tools for entanglement classification are usually difficult to implement in realistic experimental settings and are not universal due to the large number of entanglement classes [Guhne_2009]. The problem of assigning a multipartite state to some entanglement class is hard even for three qubits [Bengtsson_2006, Sabin_2008, e20100745, DeChiara_2018, PhysRevA.92.042329, Srivastava_2024, Xie2024-du], and is even more complicated for four qubits, where the number of entanglement classes is already infinite [Verstraete_2002, 10.1063/1.3511477, 10.1063/1.4946895]. Hence the variety of multipartite entanglement measures, such as based on local [PhysRevA.95.022301] or on the collective measurements [PhysRevA.69.052327]. Research has focused on limiting the experimental effort [PhysRevLett.117.210504] and recently a method called entanglement detection length has been proposed to witness many-body entanglement with a minimal number of observables [shi2024entanglement]. A technique has been developed for detecting multipartite entanglement based on a constant number of measurements, independent of the number of qubits, suitable for the Greenberger-Horne-Zeilinger (GHZ) states, cluster- or Dicke states [PhysRevLett.112.155304, PhysRevLett.103.020504]. The urgency for a versatile entanglement certification method is underlined by the fact that the Dicke states have been created in experiments with few qubits in different configurations such as photonic systems [PhysRevLett.103.020504, PhysRevA.83.013821, PhysRevLett.103.020503, Zhao:15, Wang2016, Mu:20], ultracold atoms [PhysRevLett.112.155304, doi:10.1126/science.1208798, Hamley2012, doi:10.1073/pnas.1715105115] and quantum circuits [9951196, Narisada2023]. A Hilbert-Schmidt distance between a given state and the closest separable state is another important entanglement witness [PhysRevA.102.012409]. The other methods to distinguish entanglement of quantum states use local unitary optimisations  [Rudnicki_2011, Rudnicki_2012]. The next measures are entanglement of formation [10.5555/2011326.2011329, PhysRevLett.80.2245, Luo2008, Patrick_M_Hayden_2001, Adam_W_Majewski_2002, FEI2003333, Zhao_Hui_2007, Gao2008, PhysRevA.102.042408], quantum discord [https://doi.org/10.1002/andp.200051211-1204, L_Henderson_2001, PhysRevLett.88.017901, PhysRevA.77.042303, PhysRevA.79.042325, PhysRevLett.105.030501, PhysRevA.88.014302, PhysRevLett.112.210401, Bera_2018], and concurrence [PhysRevLett.78.5022, 10.1063/1.2795840, PhysRevA.105.052441, Bhaskara2017, PhysRevA.64.042315, PhysRevA.86.062303].
@@
Another method is based on the measurement of macroscopic observables such as magnetic susceptibility [Wiesniak_2005], heat capacity [PhysRevB.78.064108], energy [brukner2004macroscopic, PhysRevA.70.062113, PhysRevA.71.010301], mean spin and spin correlations [PhysRevA.72.032309, Frydryszak2017, KUZMAK2020126579], with magnetic resonance spectroscopy [PhysRevA.100.022330, Lazarev2020], entropic correlations [PhysRevResearch.2.043152], out-of-time-order correlations [Lewis-Swan2019, PhysRevA.106.042429, PRXQuantum.2.020339] or the energy of a thermal state in equilibrium [PhysRevResearch.5.013158]. Next, there is a family of methods based on SLOCC (stochastic local operations and classical communications) and tensor properties [PhysRevA.62.062314, PhysRevA.109.032424, PhysRevLett.110.030501, 10.5555/2011734.2011739, PhysRevLett.111.110502, PhysRevA.84.062306, Zangi_2017, PhysRevLett.111.060502, Palazuelos2022genuinemultipartite]. These methods can generally be divided into subgroups [Guhne_2009]: eigenvalue criteria  [Horodecki_1996, Peres_1996, Chen_2003, Rudolph_2005, PhysRevLett.106.190502], entanglement witnesses [Guhne_2006, TERHAL2002313, PhysRevA.63.044304, Acin_2001_mixed, PhysRevLett.92.087902, Horodecki_2003, PhysRevA.77.062304, PhysRevX.8.021072, PhysRevLett.127.040401, Rico_2024], geometric and robustness criteria [Vidal_1999, PhysRevA.68.042307, 10.1063/1.1497700, doi:10.1142/S1230161222500111, 9605268, Guo_2022, quantum2010004, PhysRevLett.127.040403, PhysRevResearch.4.023059], and operational measures [Klyachko2006, Plenio_2007, PhysRevLett.115.150502]. Entanglement has been also classified with the help of mathematical tools from group theory [PhysRevA.74.022318], topology [PhysRevD.107.126005], and algebraic geometry [PhysRevResearch.2.043003]. Finally, in recent years the problem of entanglement classification in many-body systems has been approached from a machine learning point of view [PhysRevA.98.012315, Harney_2020, Harney_2021, Chen_2021, s22186767, ayachi2022general, https://doi.org/10.1002/qute.202200025, Vintskevich_2023, Giordano_2022, Asif2023, Sanavio2023, pawlowski2023identification, doi:10.1126/sciadv.add7131, palmieri2023enhancing, PhysRevA.109.022405, dawid2023modern, Wiesniak2023, urena2024entanglement].
@@ II.1 Many-body correlator
Assuming quantum-mechanical behavior of each parties of the system, i.e., $\big{|}\langle\hat{\sigma}_{+}^{(k)}\rangle\big{|}^{2}\leq 4^{-1}$, we obtain that for a separable state of $N$ qubits the following holds: ${\cal E}\leq 4^{-N}$
Based on the above considerations, to witness and classify the multi-qubit entanglement pf an $N$-qubit density matrix $\hat{\varrho}$, we use the generalized $N$-body correlator
@@ II.2 Quantum state structure characterization
For all $N$-separable states, $k=N$, $\hat{\varrho}_{n_{1}\dots n_{k}}$, $n_{l}=1$, $l\in\{1\dots N\}$, the correlator $\mathcal{Q}$ takes non-positive values
@@
and violation of this inequality indicates entanglement in the system. On the other hand, the maximal value is obtained for a non-2-separable state $\hat{\varrho}_{N}=|\psi_{\rm ghz}\rangle\langle\psi_{\rm ghz}|$, being a pure Greenberg-Horne-Zeilinger state $|\psi_{\rm ghz}\rangle=\frac{1}{\sqrt{2}}\left(|1\rangle^{\otimes N}+|0\rangle^{\otimes N}\right)$,
@@
For states being a product of a single qubit and entangled state of the remaining $N-1$ qubits, $\hat{\varrho}_{1,N-1}$, the correlator is bounded by
@@
Next, for even number of qubits $N=2M$, when $\hat{\varrho}$ is a product of $N/2$ pairs of entangled qubits, $\hat{\varrho}_{n_{1}\dots n_{k}}$, $n_{l}=2$, $l=\{1\dots N/2\}$, the correlator is bounded from above as
@@
As an examples, let us consider $N=5$ and $N=6$. For $N=5$ qubits, the density operator $\hat{\varrho}$ can take the form $\hat{\varrho}_{1,4}$, $\hat{\varrho}_{2,3}$ or $\hat{\varrho}_{1,2,2}$. The respective value of the correlator are bounded by: $\mathcal{Q}_{1,4}\leqslant 3$ and its violation certifies genuine entanglement $d=5$, $\mathcal{Q}_{2,3}\leq 3$ and its violation certifies entanglement depth $d_{e}=4$; $\mathcal{Q}_{1,2,2}\leqslant 2$ and its violation certifies entanglement depth $d_{e}=3$. For $N=6$, the entangled state $\hat{\varrho}$ can have the following structure: $\hat{\varrho}_{1,5}$, $\hat{\varrho}_{2,4}$, $\hat{\varrho}_{3,3}$, $\hat{\varrho}_{2,2,2}$. The corresponding bounds are: $\mathcal{Q}_{1,5}\leqslant 4$, $\mathcal{Q}_{2,4}\leqslant 4$, $\mathcal{Q}_{3,3}\leqslant 4$, and $\mathcal{Q}_{2,2,2}\leqslant 3$. These observations, adapted to the $N=3,4,5,6$ are summarized in Tab. 1.
@@ II.3 Extraction of $\mathcal{Q}$
This observation will be relevant in the forthcoming sections of this work. Importantly for its entanglement measure interpretation, $\mathcal{Q}$ is convex in the input states, see App. A.
@@
where, for instance, $|0_{x}\rangle$ denotes the eigenstate of $\hat{\sigma}_{x}$ with the spin pointing up. Only the correct choice of the three axes: all operators rising the spin in the $z$-direction in the former, and in the $x$, $y$ and $z$ directions in the latter case, allow to maximize the value of $\mathcal{Q}$. In what follows, we always display results for the correlator, optimized over the choice of the separable basis. What is crucial from our perspective of classification of entanglement, the value of correlations is invariant under local unitary operations. Hence the proposed method for entanglement classification respects SLOCC entanglement classes, see App. B.
@@ III Results
The number of entanglement classes grows rapidly with the number of parties and the local dimension. For pure states of two qubits, there are only two entanglement classes, with a Bell state as an extremal representative of one and separable states forming the other. Increasing the number of qubits to three yields two disjoint genuinely multipartite entanglement classes, together forming six different classes. Interestingly, the growth of the number of classes is not even exponential; already for four qubits, there are infinitely many entanglement classes, which can be divided into nine [Verstraete_2002, Zangi_2017, Vintskevich_2023] or forty-nine groups [Li_2009, Giordano_2022].
For mixed states $\hat{\varrho}$, the situation is even more complicated. The usual method of assigning $\hat{\varrho}$ to some class requires the knowledge if it can be written as a convex sum of pure states from a given SLOCC class. Notwithstanding, for a complete characterization of all mixed states, one requires also to extend it to pure states of lower entanglement, that is, non-genuine one [Acin_2001_mixed].
@@ III.1.1 Pure states
In three-qubit system, there are two types of entanglement: genuine tripartite entanglement and the one in which one qubit is separated from the rest. In the genuine entanglement case, there are two inequivalent (under SLOCC) types of states: GHZ and W, defined in App. C  [^1] The former class encompasses the majority of states in the Haar-measure sense; a state, drawn at random, will belong to the GHZ class with probability 1.. One of the key differences between both genuinely entangled states GHZ and W is their tensor rank, see App. C. Therefore, in total there are six classes: two genuine, three two-partite entangled (AB-C, AC-B, BC-A) and one separable class.
@@
Figure 2: Probability distribution of $\mathcal{Q}$ for three-qubit pure states from SLOCC classes: bi-separable-, W-, and GHZ-type (from top to bottom). The green dashed lines correspond to the maximal values of $\mathcal{Q}$ for each class. Red dashed lines presents numerically obtained probability distribution function.
@@
To assign a state to one of two given classes one can process it through the random SLOCC many times, calculating the correlator for each case. This procedure will give one of the two shapes from Fig. 2. The degree of distinguishability can be controlled with the number of samples. This protocol can be thought of as a black-box device, with random SLOCC class states on the input that, as an output, produces the correct classification provided sufficiently many samples.
@@ III.1.2 Mixed states
Figure 3: The value of the correlator $\mathcal{Q}$ for extremal representatives from each of three-qubit entanglement class, upon which we act with dephasing (dashed lines) and depolarizing (solid lines) noise. The red lines correspond to the noisy GHZ states, the blue one to W states, and the green line depicts the decay of correlations for two-qubit Bell states.
@@ III.1.3 Application: dynamical generation of entanglement
Figure 4: Dynamical generation of the many-body entanglement in spin squeezing protocols in three-qubit system, i.e., $\mathcal{Q}$ versus the time $t$. The red solid line corresponds to the one-axis twisting protocol (OAT), while the blue dashed line corresponds to the two-axis counter twisting protocol (TACT). Both protocols generate genuine entanglement at the same critical time $t_{c}$, when $\mathcal{Q}(t_{c})$ exceeds unity. The OAT protocol generates the most entangled state, namely the GHZ state, for which the correlator takes the maximal value $\mathcal{Q}=2$.

### v2

@@ I Introduction
One of the important consequences of the quantum description of physical systems is the notion of entanglement [horodecki2009entanglement] which is the necessary resource for the Einstein-Podolsky-Rosen steering [uola2020steering] and the Bell correlations [bell, brunner2014bell, bell_rmp]. None of these correlations has a classical counterpart. The highlighted role of entanglement requires a systematic characterisation and classification of quantum states in terms of some entanglement measure [PhysRevLett.78.2275, Horodecki2001, Plenio_2007]. Nevertheless, the tools for entanglement classification are usually difficult to implement in realistic experimental settings and are not universal due to the large number of entanglement classes [Guhne_2009]. The problem of assigning a multipartite state to some entanglement class is hard even for three qubits [Bengtsson_2006, Sabin_2008, e20100745, DeChiara_2018, PhysRevA.92.042329, Srivastava_2024, Xie2024-du], and is even more complicated for four qubits, where the number of entanglement classes is already infinite [Verstraete_2002, 10.1063/1.3511477, 10.1063/1.4946895]. Hence the variety of multipartite entanglement measures, such as based on local [PhysRevA.95.022301] or on the collective measurements [PhysRevA.69.052327]. Research has focused on limiting the experimental effort [PhysRevLett.117.210504] and recently a method called entanglement detection length has been proposed to witness many-body entanglement with a minimal number of observables [shi2024entanglement]. A technique has been developed for detecting multipartite entanglement based on a constant number of measurements, independent of the number of qubits, suitable for the Greenberger-Horne-Zeilinger (GHZ) states, cluster- or Dicke states [PhysRevLett.112.155304, PhysRevLett.103.020504]. The urgency for a versatile entanglement certification method is underlined by the fact that the Dicke states have been created in experiments with few qubits in different configurations such as photonic systems [PhysRevLett.103.020504, PhysRevA.83.013821, PhysRevLett.103.020503, Zhao:15, Wang2016, Mu:20], ultracold atoms [PhysRevLett.112.155304, doi:10.1126/science.1208798, Hamley2012, doi:10.1073/pnas.1715105115] and quantum circuits [9951196, Narisada2023]. A Hilbert-Schmidt distance between a given state and the closest separable state is another important entanglement witness [PhysRevA.102.012409]. The other methods to distinguish entanglement of quantum states use local unitary optimisations  [Rudnicki_2011, Rudnicki_2012]. The next measures are entanglement of formation [10.5555/2011326.2011329, PhysRevLett.80.2245, Luo2008, Patrick_M_Hayden_2001, Adam_W_Majewski_2002, FEI2003333, Zhao_Hui_2007, Gao2008, PhysRevA.102.042408], quantum discord [https://doi.org/10.1002/andp.200051211-1204, L_Henderson_2001, PhysRevLett.88.017901, PhysRevA.77.042303, PhysRevA.79.042325, PhysRevLett.105.030501, PhysRevA.88.014302, PhysRevLett.112.210401, Bera_2018], and concurrence [PhysRevLett.78.5022, 10.1063/1.2795840, PhysRevA.105.052441, Bhaskara2017, PhysRevA.64.042315, PhysRevA.86.062303, Mintert_2005, Ohnemus_2023].
@@
Another method is based on the measurement of macroscopic observables such as magnetic susceptibility [Wiesniak_2005], heat capacity [PhysRevB.78.064108], energy [brukner2004macroscopic, PhysRevA.70.062113, PhysRevA.71.010301], mean spin and spin correlations [PhysRevA.72.032309, Frydryszak2017, KUZMAK2020126579], with magnetic resonance spectroscopy [PhysRevA.100.022330, Lazarev2020], entropic correlations [PhysRevResearch.2.043152], out-of-time-order correlations [Lewis-Swan2019, PhysRevA.106.042429, PRXQuantum.2.020339] or the energy of a thermal state in equilibrium [PhysRevResearch.5.013158]. Next, there is a family of methods based on SLOCC (stochastic local operations and classical communications) and tensor properties [PhysRevA.62.062314, PhysRevA.109.032424, PhysRevLett.110.030501, 10.5555/2011734.2011739, PhysRevLett.111.110502, PhysRevA.84.062306, Zangi_2017, PhysRevLett.111.060502, Palazuelos2022genuinemultipartite]. These methods can generally be divided into subgroups [Guhne_2009]: eigenvalue criteria  [Horodecki_1996, Peres_1996, Chen_2003, Rudolph_2005, PhysRevLett.106.190502], entanglement witnesses [Guhne_2006, TERHAL2002313, PhysRevA.63.044304, Acin_2001_mixed, PhysRevLett.92.087902, Horodecki_2003, PhysRevA.77.062304, PhysRevX.8.021072, PhysRevLett.127.040401, Rico_2024], geometric and robustness criteria [Vidal_1999, PhysRevA.68.042307, 10.1063/1.1497700, doi:10.1142/S1230161222500111, 9605268, Guo_2022, quantum2010004, PhysRevLett.127.040403, PhysRevResearch.4.023059], and operational measures [Klyachko2006, Plenio_2007, PhysRevLett.115.150502]. Entanglement has been also classified with the help of mathematical tools from group theory [PhysRevA.74.022318], topology [PhysRevD.107.126005], and algebraic geometry [PhysRevResearch.2.043003]. Finally, in recent years the problem of entanglement classification in many-body systems has been approached from a machine learning point of view [PhysRevA.98.012315, Harney_2020, Harney_2021, Chen_2021, s22186767, ayachi2022general, https://doi.org/10.1002/qute.202200025, Vintskevich_2023, Giordano_2022, Asif2023, Sanavio2023, pawlowski2023identification, doi:10.1126/sciadv.add7131, palmieri2023enhancing, PhysRevA.109.022405, dawid2023modern, Wiesniak2023, urena2024entanglement, brunner2024, fuchs2024].
@@ II.1 Many-body correlator
Assuming quantum-mechanical behavior of each parties of the system, i.e., $\big{|}\langle\hat{\sigma}_{+}^{(k)}\rangle\big{|}^{2}\leq 4^{-1}$, we obtain that for a separable state of $N$ qubits the following holds
${\cal E}\leqslant 4^{-N}.$ (6)
Based on the above considerations, to witness and classify the multi-qubit entanglement of an $N$-qubit density matrix $\hat{\varrho}$, we use the generalized $N$-body correlator
@@ II.2 Quantum state structure characterization
For all $N$-separable states, $k=N$, $\hat{\varrho}_{n_{1}\dots n_{k}}$, $n_{l}=1$, $l\in\{1\dots N\}$, according to Eqs (6) and (7), the correlator $\mathcal{Q}$ takes non-positive values
@@
and violation of this inequality indicates entanglement in the system. On the other hand, the maximal value is obtained for a non-2-separable state $\hat{\varrho}_{N}=|\psi_{\rm ghz}\rangle\langle\psi_{\rm ghz}|$, being a pure Greenberg-Horne-Zeilinger state, which according to Eq. (5) gives $\mathcal{E}=1/4$, hence $|\psi_{\rm ghz}\rangle=\frac{1}{\sqrt{2}}\left(|1\rangle^{\otimes N}+|0\rangle^{\otimes N}\right)$,
@@
For states being a product of a single qubit and entangled state of the remaining $N-1$ qubits, $\hat{\varrho}_{1,N-1}$, the correlator is maximal, when the $N-1$ qubits form a GHZ state, hence again using Eq. (5), we obtain
@@
This procedure can be applied to other configurations, for instance for even number of qubits $N=2M$, when $\hat{\varrho}$ is a product of $N/2$ pairs of entangled qubits, $\hat{\varrho}_{n_{1}\dots n_{k}}$, $n_{l}=2$, $l=\{1\dots N/2\}$, the correlator is bounded from above as
@@
As examples, let us consider $N=5$ and $N=6$. For $N=5$ qubits, the density operator $\hat{\varrho}$ can take the form $\hat{\varrho}_{1,4}$, $\hat{\varrho}_{2,3}$ or $\hat{\varrho}_{1,2,2}$. The respective value of the correlator are bounded by: $\mathcal{Q}_{1,4}\leqslant 3$ and its violation certifies genuine entanglement $d_{e}=5$, $\mathcal{Q}_{2,3}\leq 3$ and its violation certifies entanglement depth $d_{e}=4$; $\mathcal{Q}_{1,2,2}\leqslant 2$ and its violation certifies entanglement depth $d_{e}=3$. For $N=6$, the entangled state $\hat{\varrho}$ can have the following structure: $\hat{\varrho}_{1,5}$, $\hat{\varrho}_{2,4}$, $\hat{\varrho}_{3,3}$, $\hat{\varrho}_{2,2,2}$. The corresponding bounds are: $\mathcal{Q}_{1,5}\leqslant 4$, $\mathcal{Q}_{2,4}\leqslant 4$, $\mathcal{Q}_{3,3}\leqslant 4$, and $\mathcal{Q}_{2,2,2}\leqslant 3$. These observations, adapted to the $N=3,4,5,6$ are summarized in Tab. 1.
@@ II.3 Extraction of $\mathcal{Q}$
This observation will be relevant in the forthcoming sections of this work. Importantly for its entanglement measure interpretation, $\mathcal{Q}$ is convex in the input states, see Appendix A.
@@
where, for instance, $|0_{x}\rangle$ denotes the eigenstate of $\hat{\sigma}_{x}$ with the spin pointing up. Only the correct choice of the three axes allows for maximizing the value of the GHZ coherence, i.e. all operators rising the spin in the $z$-direction in the former, and in the $x$, $y$ and $z$ directions in the latter case, allow to maximize the value of $\mathcal{Q}$. In what follows, we always display results for the correlator, optimized over the choice of the separable basis. What is crucial from our perspective of classification of entanglement, the value of correlations is invariant under local unitary operations. Hence the proposed method for entanglement classification respects SLOCC entanglement classes 111 The two pure multipartite quantum states $|\psi\rangle$ and $|\phi\rangle$ belong to the same SLOCC entanglement class if and only if it is possible to convert $|\psi\rangle\mapsto|\phi\rangle$ and $|\phi\rangle\mapsto|\psi\rangle$ with non-zero probabilities using local operations and classical communication, i.e., if there exist a set of invertible operators $\hat{O}_{k}$ (so that the following condition is symmetric) such that $|\phi\rangle=\bigotimes_{k=1}^{N}\hat{O}_{k}|\psi\rangle$, where $N$ is the number of constituents..
@@ III Results
The number of entanglement classes grows rapidly with the number of parties and the local dimension. For pure states of two qubits, there are only two entanglement classes, with a Bell state as an extremal representative of one and separable states forming the other. Increasing the number of qubits to three yields two disjoint genuinely multipartite entanglement classes, together forming six different classes. Interestingly, the growth of the number of classes is not even exponential; already for four qubits, there are infinitely many entanglement classes, which can be divided into nine [Verstraete_2002, Zangi_2017, Vintskevich_2023] or forty-nine groups [Li_2009, Giordano_2022]. The entanglement classification of mixed states, i.e. statistical mixtures of different pure states, is even harder, as each pure state can belong to different entanglement class [Acin_2001_mixed].
@@ III.1.1 Pure states
In three-qubit system, there are two types of entanglement: genuine tripartite entanglement and the one in which one qubit is separated from the rest. In the genuine entanglement case, there are two inequivalent (under SLOCC) types of states: GHZ and W, defined in Appendix B  [^2] The former class encompasses the majority of states in the Haar-measure sense; a state, drawn at random, will belong to the GHZ class with probability 1.. One of the key differences between both genuinely entangled states GHZ and W is their tensor rank, see Appendix B. Therefore, in total there are six classes: two genuine, three two-partite entangled (AB-C, AC-B, BC-A) and one separable class.
@@
Figure 2: Probability distribution of $\mathcal{Q}$ for three-qubit pure states from SLOCC classes: bi-separable-, W-, and GHZ-type (from top to bottom). The green dashed lines correspond to the maximal values of $\mathcal{Q}$ for each class. Red dashed lines presents numerically obtained probability distribution function. Dimensionless units.
@@
To assign a state to one of three given classes one can process it through the random SLOCC many times, calculating the correlator for each case. This procedure will give one of the three shapes from Fig. 2. The degree of distinguishability can be controlled with the number of samples. This protocol can be thought of as a black-box device, with random SLOCC class states on the input that, as an output, produces the correct classification provided sufficiently many samples.
@@ III.1.2 Mixed states
Figure 3: The value of the correlator $\mathcal{Q}$ for extremal representatives from each of three-qubit entanglement class, upon which we act with dephasing (dashed lines) and depolarizing (solid lines) noise. The red lines correspond to the noisy GHZ states, the blue one to W states, and the green line depicts the decay of correlations for two-qubit Bell states. Dimensionless units.
@@ III.1.3 Application: dynamical generation of entanglement
Figure 4: Dynamical generation of the many-body entanglement in spin squeezing protocols in three-qubit system, i.e., $\mathcal{Q}$ versus the time $t$. The red solid line corresponds to the one-axis twisting protocol (OAT), while the blue dashed line corresponds to the two-axis counter twisting protocol (TACT). Both protocols generate genuine entanglement at the same critical time $t_{c}$, when $\mathcal{Q}(t_{c})$ exceeds unity. The OAT protocol generates the most entangled state, namely the GHZ state, for which the correlator takes the maximal value $\mathcal{Q}=2$. Dimensionless units.